OpenStudy (anonymous):

solve for the equation. (1/3)x + 3=-(7/2)x-(3/2) Can I just get the answer for x? I want to see if my answer is correct....

OpenStudy (anonymous):

((1)/(3))x+3=-((7)/(2))*x-((3)/(2)) Multiply -1 by x to get -x. ((1)/(3))x+3=-x((7)/(2))-((3)/(2)) Multiply -x by each term inside the parentheses. ((1)/(3))x+3=-(7x)/(2)-((3)/(2)) Multiply -1 by the (3)/(2) inside the parentheses. ((1)/(3))x+3=-(7x)/(2)-(3)/(2) Combine the numerators of all expressions that have common denominators. ((1)/(3))x+3=(-7x-3)/(2) Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. ((1)/(3))x=-3+(-7x-3)/(2) Simplify the right-hand side of the equation. ((1)/(3))x=(-7x-9)/(2) Multiply each term in the equation by 2. ((1)/(3))x*2=(-7x-9)/(2)*2 Simplify the left-hand side of the equation by multiplying out all the terms. 2x((1)/(3))=(-7x-9)/(2)*2 Simplify the right-hand side of the equation by simplifying each term. 2x((1)/(3))=-7x-9 Multiply 2x by each term inside the parentheses. (2x)/(3)=-7x-9 Multiply each term in the equation by 3. (2x)/(3)*3=-7x*3-9*3 Simplify the left-hand side of the equation by canceling the common factors. 2x=-7x*3-9*3 Simplify the right-hand side of the equation by multiplying out all the terms. 2x=-21x-27 Since -21x contains the variable to solve for, move it to the left-hand side of the equation by adding 21x to both sides. 2x+21x=-27 Since 2x and 21x are like terms, add 21x to 2x to get 23x. 23x=-27 Divide each term in the equation by 23. (23x)/(23)=-(27)/(23) Simplify the left-hand side of the equation by canceling the common factors. x=-(27)/(23)

OpenStudy (anonymous):

so, x=-(27)/(23)

OpenStudy (anonymous):

Do you understand the concept now?

OpenStudy (anonymous):

YES! Thank you! :)